OpenStudy (anonymous):

what is the derivative of 2sinx and 2cos2x? i need to find the rules in finding derivative of such functions.

hartnn (hartnn):

know chain rule ?

hartnn (hartnn):

for 2sin x its just 2 cos x because the constant '2' can be taken out of derivative.

OpenStudy (anonymous):

product rule shows up somewhere. i know 2sinx = 2cosx and 2cos2x = -4sin2x

OpenStudy (anonymous):

*** i mean chain...rule.

hartnn (hartnn):

yes, to differentiate cos 2x ,you need chain rule, know what is it ?

OpenStudy (anonymous):

i know what it is but i don't know how to apply it to cos2x. i can work it on something like f(x) = (3x+4)^5 but not cos2x

hartnn (hartnn):

d/dx [cos 2x] = [-sin 2x] d/dx [2x] there's an inner function, and there's an outer function. d/dx composite = d/dx outer * d/dx inner here, outer = cos , inner =2x

hartnn (hartnn):

like in f(x) = (3x+4)^5 outer = x^5 inner = 3x+4

hartnn (hartnn):

so, you multiply 5(3x+4)^4 by d/dx(3x+4), right ? similarly,here you multiply by d/dx [2x] making any sense ? :P

OpenStudy (anonymous):

so i can say 2tan2x = 4sec^2x and tan2x = 2sec^2x? right?

hartnn (hartnn):

yes! correct :)

OpenStudy (anonymous):

thanks :)